OpenStudy (anonymous):

The sets of numbers 3, 4, 5 and 8, 15, 17 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples!

OpenStudy (anonymous):



OpenStudy (anonymous):

so 3^2+4^2=5^2

OpenStudy (anonymous):

Do you know the next step?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok so simplify each number. 3^2=9 4^2=16 5^2=25

OpenStudy (anonymous):

so now it looks like this 9+16=25 or 25=25 this is correct becasue a^2+b^2=c^2

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

are you still there?

OpenStudy (anonymous):

well... just do the same thing on the next one and tag me if you need help